### v1

@@ Abstract
Incomplete factorizations have long been popular general-purpose algebraic preconditioners for solving large sparse linear systems of equations. Guaranteeing the factorization is breakdown free while computing a high quality preconditioner is challenging. A resurgence of interest in using low precision arithmetic makes the search for robustness more urgent and tougher. In this paper, we focus on symmetric positive definite problems and explore a number of approaches: a look-ahead strategy to anticipate break down as early as possible, the use of global shifts, and a modification of an idea developed in the field of numerical optimization for the complete Cholesky factorization of dense matrices. Our numerical simulations target highly ill-conditioned sparse linear systems with the goal of computing the factors in half precision arithmetic and then achieving double precision accuracy using mixed precision refinement.
@@ 1 Introduction
Our interest is in solving large-scale symmetric positive definite (SPD) linear systems of equations $Ax=b$. Incomplete Cholesky (IC) factorizations of the form $A\approx LL^{T}$, where the factor $L$ is a sparse lower triangular matrix, have long been important and well-used algebraic preconditioners for use with iterative solvers. While they are general purpose and, when compared to sparse direct solvers, require modest computational resources, they do have potential drawbacks. Their effectiveness can be highly application dependent and, although significant effort has gone into developing a strong theoretical background, most results are limited to model problems. To be useful in practice, their computation and application must be efficient and robust.
@@
Traditionally, matrix factorizations have most often been computed using double precision floating-point arithmetic, which nowadays corresponds to a 64-bit floating-point number format. However, half precision arithmetic is being increasingly supported by modern hardware and because it can offer speed benefits while using less energy and memory, there has been significant interest in recent years in its potential use in numerical linear algebra; see the comprehensive review [1] and references therein. For linear systems, one strategy that has received attention is GMRES-IR [2]. The idea is to compute the matrix factors in low precision arithmetic and then employ them as preconditioners for GMRES within mixed precision iterative refinement (see also [3, 4, 1] and, for SPD systems, [5]). Using incomplete factors potentially enables much larger problems to be solved (normally at the cost of more iterations to achieve the requested accuracy). In an initial study [6], we explored this approach, focusing on the safe detection of overflows that can occur when computing matrix factors using half precision arithmetic. In the current paper, we again consider half precision IC factorizations and their potential for use in solving ill-conditioned sparse SPD linear systems to double precision accuracy. Our interest lies in exploring strategies to prevent breakdown during the factorization. Importantly, these must be robust, inexpensive and not cause serious degradation to the preconditioner quality.
@@
When using any precision, breakdown can occur during an incomplete factorization but it is more likely when using low precision arithmetic. This is partly because the initial matrix $A$ must be “squeezed” into half precision , and this may mean the resulting matrix may not be (sufficiently) positive definite [7]. But, in addition, overflows outside the narrow range of possible numerical values are an ever-present danger. Unless such overflows are correctly anticipated, the computation will crash. Here we consider and compare the performance in half precision arithmetic of a number of strategies to limit the likelihood of breakdown: look ahead that checks the diagonal entries of the partially factorized matrix at each major step of the factorization, the use of global shifts, and an approach based on locally modifying the factorization. The latter was originally used to modify approximate (dense) Hessian matrices in the field of numerical optimization. We also demonstrate that, even if double precision is used throughout the computation, it is essential to take action to prevent growth in the factor entries.
@@
The paper is organised as follows. Section 2 looks at the different stages at which breakdown can occur within an incomplete factorization. In Section 3, we present a number of ways to prevent and handle breakdown. Numerical results for a range of highly ill-conditioned linear systems coming from practical applications are presented in Section 4. Finally, in Section 5, our findings and conclusions are summarised.
Terminology. We use the term high precision for precision formats that provide high accuracy at the cost of a larger memory volume (in terms of bits) and low precision to refer to precision formats that compose of fewer bits (smaller memory volume) and provide low(er) accuracy. We are referring to IEEE double precision (64-bit) when using the term high precision (denoted by fp64) and the 1985 IEEE standard 754 half precision (16-bit) when using the term low precision (denoted by fp16). bfloat16 is another form of half-precision arithmetic that was introduced by Google in its tensor processing units and formalized by Intel; we do not use it in this paper. Table 1 summarises the parameters for different precision arithmetic.
[…]
@@ 2 Possible breakdowns within IC factorizations
A myriad of approaches for computing incomplete factorizations of sparse matrices have been developed, modified and refined over many years. Some combine the factorization with an initial step that discards small entries in $A$ (sparsification). For details of possible variants, we recommend [8, 9], while a comprehensive discussion of early strategies can be found in [10]; see also [11] for a short history and the recent monograph [12] for a general overview and skeleton algorithms.
Algorithm 1 outlines a basic (right-looking) incomplete Cholesky (IC) factorization of a sparse SPD matrix $A=\{a_{ij}\}$. It assumes a target sparsity pattern ${\mathcal{S}}\{L\}$ for the incomplete factor $L=\{l_{ij}\}$ is provided, where
@@
Modifications can be made to incorporate threshold dropping strategies and to determine ${\mathcal{S}}\{L\}$ as the method proceeds. At each major step $k$, outer product updates are applied to the part of the matrix that has yet to be factored (Lines 7–11).
@@
1:Initialize $l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
2:for $k=1:n$ donormal-$\triangleright$ Start of $k$-th major step
@@
6:     end fornormal-$\triangleright$ Column $k$ of $L$ has been computed
7:     for $j=k+1:n$ such that $(j,k)\in{\mathcal{S}}\{L\}$ do
8:         for $i=j:n$ such that $(i,j)\in{\mathcal{S}}\{L\}$  do
@@
Unfortunately, unlike a complete Cholesky factorization, there is no guarantee in general that an IC algorithm will not break down or it may exhibit large growth in the size of the factor entries (even when using double precision arithmetic). This is illustrated by the following well-conditioned SPD matrix in which $\delta>0$ is small
@@
An $IC(0)$ factorization limits entries in the incomplete factor to positions corresponding to entries in $A$, that is, ${\mathcal{S}}\{L\}={\mathcal{S}}\{A\}$. Choosing $\delta\ll 1$ and $c=1$ results in no growth in the entries of the $IC(0)$ factor and no breakdown. However, if $c=0$ and entries (2,4) and (4,2) are removed from ${\mathcal{S}}\{A\}$ then the $IC(0)$ factor becomes
@@
There are three places in Algorithm 1 where breakdown can potentially occur. Following [6], we refer to these as B1, B2, and B3 breakdowns.
@@
To develop a robust implementation, the possibility of overflow must be detected before it occurs (otherwise, the code will simply crash). Importantly, tests for breakdown need to use operations that cannot themselves overflow. An operation is said to be safe in the precision being used if it cannot overflow. To safely detect B1 breakdown it is sufficient to check at Line 3 of Algorithm 1 that $l_{kk}\geq\tau_{u}$, where the threshold parameter satisfies $\tau_{u}>1/x_{max}$. Typical values are $\tau_{u}=10^{-5}$ for half precision factorizations and $\tau_{u}=10^{-20}$ for double precision [6]; these are used in our reported experiments (Section 4). B2 breakdown can happen at Line 5. Let $d$ be equal to $l_{kk}$ (set at Line 3) and let $l_{max}$ denote the entry below the diagonal in column $k$ of largest absolute value, that is,
$l_{max}=\max_{i>k}\,\{|l_{ik}|:(i,k)\in{\mathcal{S}}\{L\}\}.$ (2.1)
If $l_{max}\leq x_{max}$ and $1\leq d\leq x_{max}$ then $d\geq l_{max}/x_{max}$ and it is safe to compute $l_{max}/d$ (and thus safe to scale column $k$). B3 breakdown can occur at Line 9. We give an algorithm for safely detecting B3 breakdown in [6].
Note that although we are focusing on SPD problems and IC factorizations, B1, B2 and B3 breakdowns are also possible during complete or incomplete factorizations of general sparse matrices. Indeed, for non SPD problems, B2 breakdowns in particular are more likely to occur. To illustrate how B2 breakdown can happen, consider the following nonsymmetric matrix, which has some large off-diagonal entries
@@
$A=LU=\begin{pmatrix}1&\cr-2/3&1\cr 0&-1.2&1\cr 2/3&0.8&-2/3&1\cr 1000/3&1000&5000&-400000&1\cr\end{pmatrix}\begin{pmatrix}3&-2&2&0&2\cr&5/3&-2&4/3&4/3\cr&&0.6&0.6&0.6\cr&&&0.1&0.1\cr&&&&-900\end{pmatrix}.$
When using fp16 arithmetic, B2 breakdown occurs on major step 4 because the $(5,4)$ entry overflows (-4000 is divided by 0.01).
@@ 3 Preventing and handling breakdown in IC factorizations
While the use of safe tests allows action to be taken before breakdown occurs, we want to reduce the likelihood of breakdown (and hence limit the effects of modifications that are needed to prevent breakdowns). Breakdown is much more likely to happen when using low precision arithmetic because of the greater likelihood of overflows occurring.
3.1 Preventing breakdown by prescaling
For both direct and iterative methods for solving systems of equations it is often beneficial to prescale the matrix, that is, to determine diagonal matrices $S_{r}$ and $S_{l}$ (with $S_{r}=S_{l}$ in the symmetric case) such that the scaled matrix $\widehat{A}=S_{r}^{-1}AS_{l}^{-1}$ is “nicer” than the original $A$. By nicer, we mean that, compared with solving $Ax=b$, it is easier to solve the system $\widehat{A}y=S_{r}^{-1}b$ and then set $x=S_{l}^{-1}y$. When working in fp16, scaling is essential because of the narrow range of the arithmetic (recall Table 1). Numbers of absolute value outside the interval $[x_{min}^{s},x_{max}]=[5.96\times 10^{-8},6.55\times 10^{4}]$ cannot be represented in fp16 and they underflow or overflow when converted to fp16. Moreover, to avoid the performance penalty of handling subnormal numbers, in practice numbers with small absolute values are often flushed to zero. We do this for all entries of absolute value less than $10^{-5}$ in fp16 and $10^{-20}$ in fp64. When factorizing $\widehat{A}$ in fp16 arithmetic, overflow and underflow may occur during the initial conversion (squeezing) of the matrix into fp16 as well as during the factorization. The scaling must be chosen to prevent overflows. For incomplete factorizations, numbers that underflow or are flushed to zero are not necessarily a concern because the factorization is approximate. However, the resulting sparsification may mean that the scaled and squeezed matrix is close to being indefinite.
No single approach to constructing a scaling is universally the best and sparse solvers frequently include a number of options to allow users to experiment to determine the most effective for their applications. Our experience with IC factorizations of SPD matrices is that it is normally sufficient to use simple $l_{2}$-norm scaling (that is, $S_{r}=S_{l}=D^{1/2}$ where $D_{ii}=\|Ae_{i}\|_{2}$ with $e$ the vectors of 1’s, is computed in fp64 arithmetic), resulting in the absolute values of the entries of the scaled matrix $\widehat{A}$ being at most 1. This is used in the current study (but see [13], where equilibration scaling is used and [5] where scaling by the square root of the diagonal is used).

### v4

@@ Abstract
Incomplete factorizations have long been popular general-purpose algebraic preconditioners for solving large sparse linear systems of equations. Guaranteeing the factorization is breakdown free while computing a high quality preconditioner is challenging. A resurgence of interest in using low precision arithmetic makes the search for robustness more important and more challenging. In this paper, we focus on ill-conditioned symmetric positive definite problems and explore a number of approaches for preventing and handling breakdowns: prescaling of the system matrix, a look-ahead strategy to anticipate breakdown as early as possible, the use of global shifts, and a modification of an idea developed in the field of numerical optimization for the complete Cholesky factorization of dense matrices. Our numerical simulations target highly ill-conditioned sparse linear systems with the goal of computing the factors in half precision arithmetic and then achieving double precision accuracy using mixed precision refinement. We also consider the often overlooked issue of growth in the sizes of entries in the factors that can occur when using any precision and can render the computed factors ineffective as preconditioners.
@@ 1 Introduction
Our interest is in solving large-scale symmetric positive definite (SPD) linear systems of equations $Ax=b$. Incomplete Cholesky (IC) factorizations of the form $A\approx LL^{T}$, where the factor $L$ is a sparse lower triangular matrix, have long been important and well-used algebraic preconditioners for use with iterative solvers. While they are general purpose and, when compared to sparse direct solvers, require modest computational resources, they do have drawbacks. Their effectiveness can be highly application dependent and, although significant effort has gone into developing a strong theoretical background, most results are limited to model problems. To be useful in practice, their computation and application must be efficient and robust.
@@
Traditionally, matrix factorizations have most often been computed using double precision floating-point arithmetic, which nowadays corresponds to a 64-bit floating-point number format. However, half precision arithmetic is being increasingly supported by modern hardware and because it can offer speed benefits while using less energy and memory, there has been significant interest in recent years in its use in numerical linear algebra; see the comprehensive review [1] and references therein. For linear systems, one strategy that has received attention is GMRES-IR [2]. The idea is to compute the matrix factors in low precision arithmetic and then employ them as preconditioners for GMRES within mixed precision iterative refinement (see also [3, 4, 1]). For SPD systems, GMRES can potentially be replaced by the CG (conjugate gradient) method [5]. Using low precision incomplete factors may enable much larger problems to be solved (normally at the cost of more iterations to achieve the requested accuracy). In an initial study [6], we explored this approach, focusing on the safe avoidance of overflows that can occur when computing matrix factors using low precision arithmetic.
When using any precision, breakdown can occur during an incomplete Cholesky factorization of a general SPD matrix, that is, a pivot (diagonal entry of a Schur complement) may be zero or negative where an exact Cholesky factorization would have only positive pivots, or a computation within the factorization may overflow. A review is given in the Lapack Working Note [7]. Breakdown is more likely when using low precision arithmetic. This is partly because the initial matrix $A$ must be “squeezed” into half precision, which may mean the resulting matrix is not (sufficiently) positive definite for the factorization to be successful [8]. But, in addition, overflows outside the narrow range of possible numerical values are an ever-present danger. Our interest lies in exploring strategies to prevent breakdown during sparse matrix factorizations. Importantly, these must be robust, inexpensive and not cause serious degradation to the preconditioner quality.
@@
Building on our earlier work on breakdowns within half precision sparse matrix incomplete factorizations [6], this paper makes the following contributions. Firstly, we consider and compare the performance in half precision arithmetic of a number of strategies to limit the likelihood of breakdown: (a) prescaling the matrix before it is squeezed, (b) look-ahead that checks the diagonal entries of the partially factorized matrix at each major step of the factorization, (c) the use of global shifts, and (d) an approach based on locally modifying the factorization. The latter was originally used to modify approximate (dense) Hessian matrices in the field of numerical optimization. Here we seek to apply it to sparse problems, in combination with low precision arithmetic, as a strategy that avoids the restarting needed with the use of global shifts. Secondly, we demonstrate that, even if double precision is used throughout the computation and breakdown does not occur, it is essential to take action to prevent growth in the factor entries because otherwise, the factors can be ineffective as preconditioners. Finally, we concentrate our numerical experiments on highly ill-conditioned linear systems and the challenge of recovering double accuracy in the computed solution using a preconditioner computed in low precision arithmetic. We develop Fortran software that enables us to illustrate the potential for exploiting half precision within robust approaches for tackling large-scale sparse systems.
@@
The rest of the paper is organised as follows. Section 2 looks at the different stages at which breakdown can occur within an incomplete factorization. In Section 3, we present a number of ways to prevent and handle breakdown. Numerical results for a range of highly ill-conditioned linear systems coming from practical applications are presented in Section 4. Finally, in Section 5, our findings and conclusions are summarised.
Terminology. We use high precision (denoted by fp64) to refer to IEEE double precision (64-bit) and low precision (denoted by fp16) for the 1985 IEEE standard 754 half precision (16-bit). Note that bfloat16 is another form of half precision arithmetic. It has 8 bits in the significand and (as in fp32) 8 bits in the exponent. We do not use it in this paper because our software is written in Fortran and, as far as we are aware, there are currently no Fortran compilers that support the use of bfloat16. Table 1 summarises the parameters for the precisions used in this paper.
Table 1: Parameters for fp16, fp32, and fp64 arithmetic: the number of bits in the significand (including the implicit most significant bit) and exponent, unit roundoff $u$, smallest positive (subnormal) number $x^{s}_{min}$ , smallest normalized positive number $x_{min}$, and largest finite number $x_{max}$, all given to three significant figures.
@@ 2 Possible breakdowns within IC factorizations
A myriad of approaches for computing incomplete factorizations of sparse matrices have been developed, modified and refined over many years. Some combine the factorization with an initial step that discards small entries in $A$ (sparsification). For details of possible variants, we recommend [9, 10], while a comprehensive discussion of early strategies can be found in [11]; see also [12] for a short history and the recent monograph [13] for a broad overview and skeleton algorithms. We note that significant progress in the field of preconditioning has been achieved by looking for incomplete factorizations that are breakdown-free because of the properties of the matrix $A$, for example, for M and H-matrices. A seminal paper on this is [14]; see also the summary in [15].
Algorithm 1 outlines a basic (right-looking) incomplete Cholesky (IC) factorization of a sparse SPD matrix $A=\{a_{ij}\}$ [^1] The Algorithm can be modified to compute a square-root free LDLT factorization in which $L$ has unit diagonal entries and $D$ has positive entries.. It assumes a target sparsity pattern ${\mathcal{S}}\{L\}$ for the incomplete factor $L=\{l_{ij}\}$ is provided, where
@@
The simplest case ${\mathcal{S}}\{L\}={\mathcal{S}}\{A\}$ is called an $IC(0)$ factorization. Modifications to Algorithm 1 can be made to incorporate threshold dropping strategies and to determine ${\mathcal{S}}\{L\}$ as the method proceeds. At each major step $k$, outer product updates are applied to the part of the matrix that has yet to be factored (Lines 7–11).
@@
1:$l_{ij}=a_{ij}$ for all $(i,j)\in{\mathcal{S}}\{L\}$
2:for $k=1:n$ $\triangleright$ Start of $k$-th major step
@@
6:     end $\triangleright$ Column $k$ of $L$ has been computed
7:     for $j\in\{j>k\,|\,(j,k)\in{\mathcal{S}}\{L\}\}$ do
8:         for $i\in\{i\geq k\,|\,(i,k)\in{\mathcal{S}}\{L\}\}$ do
@@
Unfortunately, unlike a complete Cholesky factorization, there is no guarantee in general that an IC algorithm will not break down or exhibit large growth in the size of the factor entries (even when using double precision arithmetic). This is illustrated by the following well-conditioned SPD matrix in which $\delta>0$ is small
@@
Choosing $\delta\ll 1$ and $c=1$ results in no growth in the entries of the $IC(0)$ factor and no breakdown. However, if $c=0$ and entries (2,4) and (4,2) are removed from ${\mathcal{S}}\{A\}$ then the $IC(0)$ factor becomes
@@
There are three places in Algorithm 1 where breakdown can occur. Following [6], we refer to these as B1, B2, and B3 breakdowns.
@@
To develop robust IC factorization implementations, breakdowns must either be avoided or they must be detected and handled by restarting the computation with revised data. We seek to avoid breakdowns but, as we cannot guarantee there will be no breakdowns, we still need to monitor for them. A recent study involving multi-precision iterative refinement used the functions offered by MATLAB to check the computed factors for Inf and/or NaN entries and took action if such entries were found [16]. This is not a practical procedure for general use. One possible strategy is to use IEEE-754 floating-point exception handling. This allows overflows to occur, the execution continues until a status flag is checked and, at this point, if overflow has been detected, restarting is initiated; see, for example, [17]. This is straightforward but requires the user to employ the correct compiler flags, which may be challenging, for instance, when a solver is interfaced from other languages. Furthermore, it is likely that non-IEEE arithmetics will gain traction in the future [18]. An alternative and potentially more flexible strategy is to incorporate explicit tests for breakdown into the factorization algorithm. In this case, for an implementation to be robust, the tests employed must only use operations that cannot themselves overflow.
An operation is said to be safe in the precision being used if it cannot overflow. To safely detect B1 breakdown it is sufficient to check at Line 3 of Algorithm 1 that $l_{kk}\geq\tau_{u}$, where the threshold parameter satisfies $\tau_{u}>1/x_{max}$. This ensures $(l_{kk})^{-1}<x_{max}$. Typical values are $\tau_{u}=10^{-5}$ for half precision factorizations and $\tau_{u}=10^{-20}$ for double precision [6]; these are used in our reported experiments (Section 4). B2 breakdown can happen at Line 5. Let $l_{kmax}$ denote the entry below the diagonal in column $k$ of largest absolute value, that is,
$l_{kmax}=\max_{i>k}\,\{|l_{ik}|:(i,k)\in{\mathcal{S}}\{L\}\}.$ (2.1)
If $l_{kmax}\leq x_{max}$ and $1\leq l_{kk}\leq x_{max}$ or $l_{kk}\geq l_{kmax}/x_{max}$ then it is safe to compute $l_{kmax}/l_{kk}$ (and thus safe to scale column $k$). B3 breakdown can occur at Line 9. We give an algorithm for safely detecting B3 breakdown in [6]. Given scalars $a,b,c$ such that $|a|,|b|,|c|\leq x_{max}$, the algorithm returns $v=a-bc$ or a flag to indicate $v$ cannot be computed safely. It does this in two stages: it first checks whether $w=bc$ can be computed safely and then whether $v=a-w$ can be computed safely.
Note that although we are focusing on SPD problems and IC factorizations, B1, B2 and B3 breakdowns are also possible during complete or incomplete factorizations of nonsymmetric sparse matrices. Indeed, for non SPD problems, B2 breakdowns (that is, overflow of one or more entries when the pivot column is divided by the pivot) in particular may be more likely to occur (although remains uncommon if the matrix is well-scaled). To demonstrate how B2 breakdown can happen, consider the following nonsymmetric matrix, which has some large off-diagonal entries
@@
$A=LU=\begin{pmatrix}1&\cr-2/3&1\cr 0&-1.2&1\cr 2/3&0.8&-2/3&1\cr 1000/3&1000&5000&-400000&1\cr\end{pmatrix}\begin{pmatrix}3&-2&2&0&2\cr&5/3&-2&4/3&4/3\cr&&0.6&0.6&0.6\cr&&&0.01&0.01\cr&&&&-900\end{pmatrix}.$
When using fp16 arithmetic, B2 breakdown occurs when performing the $4$th elimination step because the $(5,4)$ entry overflows (-4000 is divided by 0.01).
@@ 3 Preventing and handling breakdown in IC factorizations
While the use of safe tests allows action to be taken before breakdown occurs or the use of IEEE exception handling can capture breakdown, our objective is to reduce the likelihood of breakdown. This will limit the overheads involved in handling breakdowns and the effects on the quality of the computed factorizations through modifications to the data. Breakdown is much more likely to happen when using low precision arithmetic because of the greater likelihood of overflows occurring.
3.1 Avoiding breakdown by prescaling
For both direct and iterative methods for solving systems of equations it is often beneficial to prescale the matrix, that is, to determine diagonal matrices $S_{r}$ and $S_{l}$ (with $S_{r}=S_{l}$ in the symmetric case) such that the scaled matrix $\widehat{A}=S_{r}^{-1}AS_{l}^{-1}$ is “nicer” than the original $A$. By nicer, we mean that, compared with solving $Ax=b$, it is easier to solve the system $\widehat{A}y=S_{r}^{-1}b$ and then set $x=S_{l}^{-1}y$. When working in fp16 arithmetic, scaling is essential because of the narrow range of the arithmetic (recall Table 1). Numbers of absolute value outside the interval $[x_{min}^{s},x_{max}]=[5.96\times 10^{-8},6.55\times 10^{4}]$ cannot be represented in fp16 arithmetic and they underflow or overflow when converted to fp16 arithmetic. Moreover, to avoid the performance penalty of handling subnormal numbers, in practice numbers with small absolute values are often flushed to zero (that is, replaced by zero). Before factorizing $\widehat{A}$ in fp16 arithmetic, a scaling is chosen so that when converting (squeezing) the scaled matrix into fp16, overflow is avoided. In this initial squeezing of the matrix, we flush to zero all entries in the scaled matrix of absolute value less than $10^{-5}$. For incomplete factorizations, numbers that underflow or are flushed to zero are not necessarily a concern because the factorization is approximate. However, the resulting sparsification may mean that the scaled and squeezed matrix is close to being indefinite.
No single approach to constructing a scaling is universally the best and sparse solvers frequently include a number of options to allow users to experiment to determine the most effective for their applications (or to supply their own scaling). Our experience with IC factorizations of SPD matrices is that it is normally sufficient to use simple $l_{2}$-norm scaling (that is, using fp64 arithmetic, we compute $S_{r}=S_{l}=D^{1/2}$, where $d_{ii}$ is the 2-norm of row $i$ of $A$), resulting in the absolute values of the entries of the scaled matrix $\widehat{A}$ being at most 1. This is used in the current study (but see [19], where equilibration scaling is used and [5] where scaling by the square root of the diagonal is used).